From: Wolf_Charlie on Mon, Dec 30, 1996 6:37 AM Subject: Daily Report #1786 12/30/96 To: a1; a2; barnbaum; burch; c1; c2; c5; c6; h2; h8; h9; hstbb4; k4; l2; = l3; l4; Leibee_Jack; m4; m5; n1; o1; p4; p5; r1; r4; r5; r6; s7; stdesk; = t1; w1; w3; w5; w6; z1 HUBBLE SPACE TELESCOPE DAILY REPORT #1786 PERIOD COVERED: 0000Z (UTC) 12/27/96 - 0000Z (UTC) 12/30/96 Daily Status Report as of 365/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/48 6198 (Cycle 5 F/48 Long Slit Spectrophotometric = Calibration) The Faint Object Camera (f/48) was used to make observations of = the standard star LDS749B in order to perform a long slit = spectrophotometric calibration. Observations were also taken with the = star at different positions on the long slit. This will allow some = measurement of the spatial dependence of the sensitivity, and allow = tracing the distortion pattern at the different positions on the slit. = Dark and internal flat calibration images were also taken. The = observations were completed as scheduled, and no anomalies were reported. = 1.2 Completed Six Sets of WF/PC-2 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = During the second iteration, a guide star acquisition failure occurred at = 362/0827Z and defaulted to fine lock using FGS#1 only, due to a FGS#3 = scan step being exceeded. As a consequence, all the following = reacquisitions were conducted on FGS#1. HSTAR 5854 was written. All = remaining observations were completed as planned, and no further problems = were reported. 1.3 Completed Four Sets of WF/PC-2 6802 (Cosmology with The Medium = Deep Survey) The WF/PC-2 was used to make observations in parallel with the = prime science observations as part of the Medium Deep Survey. These = parallel MDS exposures have provided unique high resolution images which = clearly supplement the primary HST surveys. The observation were = executed as scheduled, and no anomalies were noted 1.4 Completed Five Sets of FOS/BL 6913 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make a = series of stability calibration measurements of the wavelength location = and the location of the spectra. All the observations were internals. = The observations were completed as planned, and no problems were = reported. 1.5 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTULIR16. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.6 Completed GHRS/2 6710 (Abundances in the Neutral Medium of the = Compact Galaxy IZW-18) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the compact galaxy IZW-18 in order to = study the abundances in its neutral medium. Previous observations with = the GHRS indicate very low levels of oxygen. A wavelength calibration = was also taken. A guide star acquisition failure occurred at 362/0827Z = and defaulted to fine lock using FGS#1 only, due to a FGS#3 scan step = being exceeded. As a consequence, all the following reacquisitions were = conducted on FGS#1. GHRS observations 52 through 62 were affected. = HSTAR 5854 was written. Otherwise, the observations were completed as = planned, and no further problems were reported. 1.7 Completed WF/PC-2 6348 (Blue Continuum Snapshots of 3CR Radio = Galaxies) The WF/PC-2 was used to make some snapshot images of the blue = continuum in radio galaxy 3C166.0. Combined with existing data, these = exposures will provide high spatial resolution information on the color = distribution within galaxies, redding due to dust, spacial indices of the = optical synchrotron jets, and colors of the active nuclei and tidal jets. = The observations were executed as scheduled, and no anomalies were = noted. 1.8 Completed GHRS/1 6707 (M-33: An Archetypal QSO Absorption Line = Galaxy) The High Resolution Spectrograph (Side One) was used to make = observations of the C IV, Si IV, and Si II spectrographic absorption = lines QSO Q0134-3254 whose light passes through the galaxy M-33. This = allows for one to probe into the interstellar medium in outer disk and = halo of galaxy M-33. The observations were completed as planned, and no = problems were noted. 1.9 Completed Four Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: = Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 1.10 Completed GHRS/1 6896 (GHRS Far UV Throughput With COSTAR Mirrors = - Cycle 6 Close-out) The High Resolution Spectrograph (Side One) was used to make = far-ultraviolet throughput calibration observations of the standard star = BD75D325 in order to study the effects caused by the COSTAR mirrors. = This is to determine if the COSTAR is the sole cause of the loss of = throughput at the short-wavelength end of the spectra. The observations = were completed as planned, and no anomalies were reported. 1.11 Completed Two Sets of GHRS/1 6900 (GHRS Echelle and G140M = Long-Term Monitor - Cycle 6) The High Resolution Spectrograph (Side One), using echelle grating = G140M, was used to make observations of standard star Mu-Col through the = large science aperture as part of the system's long term monitor and = calibration. The observations were completed as planned, and no problems = were reported. 1.12 Completed FOS/RD/WF/PC-2 6100 (Measuring the Characteristic Size = of Lyman-Alpha Forest Clouds) The Faint Object Spectrograph (Red Side) and the WF/PC-2 were both = used to make observations of quasar Q0107-025A in order to study the = characteristics and size of the intervening Lyman-Alpha forest clouds. = The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed Three Sets of WF/PC-2 7202 (The WF/PC-2 B Band Parallel = Survey: A Systematic and Synoptic Study of Galaxy Formation and Evolution = - Short Orbit Case) The WF/PC-2 was used to make observations as part of a B-Band = Survey to make a synoptic study of galaxy formation and evolution. These = observations were taken in parallel with the prime science observations. = They were completed as planned, and no anomalies were reported. 1.14 Completed GHRS/1/2 6074 (Kinematics of the Disc of TT Hydrae) The High Resolution Spectrograph (Both Sides One and Two) was used = to make spectrographic observations of the disk around binary star system = TT-Hydrae (aka, HD97528) in order to study the dynamics of its accretion = disk system. The cooler of the two stars should give rather high = accretion rates through the Roche lobe overflow. A wavelength = calibration image was also taken. The observations were completed as = scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 35 Successful : 35 2.2 FHST Updates: Scheduled : 24 Successful : 24 2.3 Operations Notes: Another quiet period was experienced by HST operations personnel = over the weekend. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The next HST Daily Report will be issued on Friday, January 3, = 1997. /DMH